Calculus II, Spring 2023 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework 01

1. Section 10.1: Problems 46, 53, 59, 63, 67, 69, 73, 81, 87, 125. Hint for problem 73: Read Appendix A.5. Hint for problem 87: Multiply by \frac{n + \sqrt{n^{2} − n}}{n + \sqrt{n^{2} − n}}. Hint for problem 125: See page 2 and page 3 in v02 of lecture 02 note to see the revised (simpler) argument on how to find N.

2. Section 10.2: Problems 31, 33, 43, 61, 65, 71, 78.